OpenStudy (anonymous):

How do I do Substitution Method?-4.5x-2y=-12.5 and 3.25x-y=-0.75

OpenStudy (anonymous):

3.25 x - y = - 0.75 solve for y. then substitute y in the other equation

OpenStudy (anonymous):

So y=3.25x+0.75 but...I'm still kinda confused

OpenStudy (anonymous):

substitute this y = 3.25 x + 0.75 into -4.5x - 2y = -12.5 -4.5x - 2(3.25x + 0.75) = - 12.5 and solve~